# Epistemic Integrity in Large Language Models

### v1

@@ Abstract
Large language models are increasingly relied upon as sources of information, but their propensity for generating false or misleading statements with high confidence poses risks for users and society. In this paper, we confront the critical problem of epistemic miscalibration — where a model’s linguistic assertiveness fails to reflect its true internal certainty. We introduce a new human-labeled dataset and a novel method for measuring the linguistic assertiveness of Large Language Models (LLMs) which cuts error rates by over 50% relative to previous benchmarks. Validated across multiple datasets, our method reveals a stark misalignment between how confidently models linguistically present information and their actual accuracy. Further human evaluations confirm the severity of this miscalibration. This evidence underscores the urgent risk of the overstated certainty LLMs hold which may mislead users on a massive scale. Our framework provides a crucial step forward in diagnosing this miscalibration, offering a path towards correcting it and more trustworthy AI across domains.
@@ 1 Introduction
Figure 1: This figure illustrates two examples with varying levels of epistemic calibration in LLM outputs. The one below is poorly calibrated. For each output, we calculate two certainty scores: internal certainty and external certainty (linguistic assertiveness). The internal certainty is computed using the method outlined by Rivera et al. (2024). To assess linguistic assertiveness, we develop a custom model, which we validate using human ratings collected through a survey.
@@
Several studies have explored the calibration of internal confidence in LLMs. For instance, Zhang et al. (2024) examine confidence calibration, proposing techniques to reduce hallucinations and enhance the model’s ability to answer known questions while avoiding unknown ones. However, they overlook the role of linguistic assertiveness and how external certainty can still lead to epistemic miscalibration even if internal confidence is addressed. Similarly, Ren et al. (2023) focus on factual knowledge and LLM behavior before and after retrieval-augmented generation (RAG). While they investigate internal confidence, they fail to frame miscalibration as an end-to-end issue involving both internal certainty and linguistic assertiveness, therefore ignoring the interplay between model predictions and how confidence is expressed linguistically.
@@
More recent studies have attempted to bridge the gap between internal confidence and linguistic assertiveness but still face considerable limitations. Although Mielke et al. (2022) explore epistemic calibration, their study uses a limited scoring scale to measure both assertiveness and confidence, which restricts the continuous assessment of LLM output. Their approach also relies on a narrow range of datasets, thereby limiting its applicability across domains. Zhou et al. (2024a) address miscalibration using epistemic markers, but their method lacks real domain grounding and fails to consider the complexity of natural language.
@@
This review of existing work on LLM calibration and confidence reveals several gaps that our research aims to address:
@@
  • Lack of Integrated Approaches: Previous studies address either internal certainty or linguistic assertiveness but rarely both simultaneously (Jiang et al., 2021). There is a need for comprehensive frameworks that integrate these aspects to ensure LLMs communicate accurately and responsibly.
@@
  • Inadequate Assertiveness Measurement: Existing methods for measuring linguistic assertiveness rely heavily on lexicon-based approaches (Pei & Jurgens, 2021) or subjective perceptions without adequate validation (Steyvers et al., 2024). These methods often lack contextual depth and fail to generalize across diverse domains.
@@
  • Limited High-Stakes Evaluation: Although some studies explore epistemic calibration, they cover a narrow range of topics and employ low-resolution measures of assertiveness, limiting their applicability in critical domains such as misinformation detection (Mielke et al., 2022).
@@
This study also presents a novel human-centered approach for developing a robust assertiveness scoring method. To ensure reliability, we train and compare several models to identify the best estimator for assertiveness based on accuracy and transferability using a new multi-domain dataset hand-coded to measure certainty. After selecting the top-performing model, we validate the results using a different human-surveyed dataset (Wang, 2017), which was coded independently by different individuals. This comprehensive methodology enables us to thoroughly assess both the objective and subjective aspects of assertiveness in language model explanations.
Our findings reveal that when the model has low internal certainty, it generates explanations that are significantly over-assertive, meaning the language used implies a higher degree of certainty than is warranted by the model’s actual confidence or accuracy. This miscalibration could lead users to misconstrue the model’s judgments as more reliable than they actually are. More precisely, our results confirm a strong correlation between GPT 4o’s assertiveness scores and human perceptions of assertiveness, but a weak correlation between human perceptions and internal certainty, and an even weaker relationship between GPT 4o model assertiveness and internal certainty. Together, these results provide the most extensive evidence of epistemic miscalibration in LLMs to date. [^1] For the code and datasets used, refer to our GitHub repository at: https://github.com/ComplexData-MILA/epistemic-integrity.
@@ 2.1.2 External Certainty
Mielke et al. (2022) examine how both humans and LLMs interpret and generate expressions of uncertainty. They find that LLMs such as GPT-4 can map uncertainty expressions to numerical equivalents similarly to humans, though they are more prone to biases rooted in their training data. However, most prior research focuses on how LLMs interpret uncertainty expressions rather than how they generate assertive or uncertain language aligned with their internal confidence scores.
@@ 2.2 Epistemic Calibration
Figure 1 presents two contrasting examples that highlight varying levels of epistemic calibration in LLM outputs. In both cases, the LLM responds to a user query, and we compute two certainty scores: Internal Certainty and External Certainty (through linguistic assertiveness). In the first example, the internal and external certainty scores are closely aligned, demonstrating high epistemic calibration. Here, the model’s linguistic assertiveness accurately reflects its internal confidence, resulting in a trustworthy output. In contrast, the second example reveals low epistemic calibration. Although the internal certainty score is relatively low, the model’s linguistic assertiveness remains high, indicating overconfidence in its response. In the following analysis, we compute internal certainty using the method outlined by Rivera et al. (2024) and use linguistic assertiveness, derived from our custom model detailed in Section 3—which is validated against human surveyed human ratings—as a proxy for external certainty quantification.
@@ 3.2 Computing internal certainty and linguistic assertiveness
To estimate the internal certainty of the LLM, we use the method outlined in Rivera et al. (2024), which discusses the state-of-the-art methods for the misinformation detection domain. The model provides an explanation for each misinformation classification and assigns an uncertainty score. The scores are calibrated on a validation set using Platt’s method (Platt et al., 1999). The robustness of this method is further validated in Appendix B.
@@ 4 Results
Figure 2 is obtained from comparing the seven different methods of assertiveness quantification, evaluating on the test set of our composite dataset. We find that GPT-4o fine-tuned with rounding (training on assertiveness scores rounded to one decimal point) achieves the highest accuracy in predicting human-annotated assertiveness scores. The margin of improvement over the approaches from the literature is very large, cutting Mean Squared Error (MSE) by more than half. To validate the transferability of these results across different domains, we conduct an ablation study by training the model on only four of the five datasets, and subsequently testing the model on the excluded one (as opposed to a standard random split in Figure 2). As shown in Table 1, GPT-4o fine tuned (with training assertiveness scores rounded to one decimal point) achieves the lowest average MSE. Thus, GPT-4o appears well-suited to capturing the linguistic nuances that contribute to perceived assertiveness, even in transfer settings, and we use it as our primary tool for measuring the assertiveness of generated explanations in the misinformation detection domain.
Figure 2: Comparison of assertiveness evaluation performance across models. The models are evaluated based on their Mean Squared Error (MSE) relative to the test set, which is a subset of our dataset. Among all models, the fine-tuned GPT-4o, trained with assertiveness scores rounded to one decimal point, achieved the lowest MSE, which suggests the highest accuracy in predicting assertiveness.
@@
Table 1: A comparison of model performance when training on four datasets and holding out the fifth for the measurements reported here. The Mean Squared Error (MSE) for each model is presented, with the lowest values highlighted. The fine-tuned GPT 4o model, using assertiveness scores rounded to one decimal place, achieves the best overall performance with an average MSE of approximately $1.0781.078$. This result indicates that the model is the most consistent and reliable for assertiveness scoring across various domains.
@@
Figure 3(a) illustrates a comparison between the probability distributions of certainty scores derived from the certainty calibration method proposed by Rivera et al. (2024) and assertiveness scores from the best-performing model shown in Figure 2 (Fine-tuned GPT, Rounded), applied to the LIAR misinformation dataset. Notably, while the certainty scores exhibit a wide variance, assertiveness scores are more concentrated toward the middle of the distribution. Additionally, Figure 3(b) reveals a low Spearman correlation (0.3) between the two sets of scores, suggesting that there is a significant misalignment between certainty and assertiveness. We provide examples of both epistemically calibrated and uncalibrated explanations with varying levels of assertiveness in Appendix F. Furthermore, in Appendix G, we test whether strong assertions of uncertainty affect calibration, finding that even when controlling for these cases, the model remains heavily skewed towards over-assertiveness.
@@ 7 Discussion and Conclusion
In this work, we introduced the problem of epistemic calibration for LLMs: ensuring that the confidence expressed in a model’s communication aligns with its underlying reliability. We argued that this normative ideal is critical for LLMs to serve as robust and responsible information sources. Through a decomposition of the problem into external and internal certainty, we developed a framework for understanding and evaluating epistemic calibration in LLMs. Using our new approach that greatly improves fidelity of assertiveness measurements compared to prior models, our empirical investigation of a state-of-the-art modeleveals significant gaps between the model’s internal confidence estimates and the assertiveness of its generated language. This miscalibration poses risks to users, who may be misled by overconfident model outputs.
@@ 10 Author Contributions
Bijean Ghafouri led the human evaluation of assertivity, took the lead writing the second (current) version of the paper, and gave input and feedback throughout the project. Shahrad Mohammadzadeh led the work towards a stronger assertiveness measurement model, and made significant contributions to the direction, writing, and other parts of the project. James Zhou had the idea of assertiveness miscalibration, created the original framing and first version of the paper, and built and ran the initial experiments. Pratheeksha Nair conducted downstream analysis of assertiveness (mis)calibration and contributed to the writing. Jacob-Junqi Tian fine-tuned Llama-3.2 and set up the framework for our composite dataset labeling. Mayank Goel expanded the assertivity evaluation experiments. Reihaneh Rabbany and Jean-François Godbout advised the project, contributing ideas and feedback. Godbout also helped design the human evaluation and contributed substantially to the writing. Kellin Pelrine supervised the project, providing guidance and feedback at all stages.

### v2

@@ Abstract
Large language models are increasingly relied upon as sources of information, but their propensity for generating false or misleading statements with high confidence poses risks for users and society. In this paper, we confront the critical problem of epistemic miscalibration — where a model’s linguistic assertiveness fails to reflect its true internal certainty. We introduce a new human-labeled dataset and a novel method for measuring the linguistic assertiveness of Large Language Models (LLMs) which cuts MSE by over 50% relative to previous benchmarks. Validated across multiple datasets, our method reveals a stark misalignment between how confidently models linguistically present information and their actual accuracy. Further human evaluations confirm the severity of this miscalibration. This evidence underscores the urgent risk of the overstated certainty LLMs hold which may mislead users on a massive scale. Our framework provides a crucial step forward in diagnosing this miscalibration, offering a path towards correcting it and more trustworthy AI across domains.
@@ 1 Introduction
Figure 1: This figure illustrates two examples with varying levels of epistemic calibration in LLM outputs. The one below is poorly calibrated. For each output, we calculate two certainty scores: internal certainty and external certainty (linguistic assertiveness). The internal certainty is computed using the method outlined by Pelrine et al. (2023). To assess linguistic assertiveness, we develop a custom model, which we validate using human ratings collected through a survey.
@@
Several studies have explored the calibration of internal confidence in LLMs. For instance, Zhang et al. (2024) examine confidence calibration, proposing techniques to reduce hallucinations and enhance the model’s ability to answer known questions while avoiding unknown ones. However, they overlook the role of linguistic assertiveness and how external certainty can still lead to epistemic miscalibration even if internal confidence is addressed. Similarly, Ren et al. (2023) focus on factual knowledge and LLM behavior before and after retrieval-augmented generation (RAG). While they investigate internal confidence, they fail to frame miscalibration as an end-to-end issue involving both internal certainty and linguistic assertiveness, therefore ignoring the interplay between model predictions and how confidence is expressed linguistically. Since this study was first conducted, emerging work has sharpened the theoretical distinction between internal confidence and linguistic assertiveness, emphasizing the potential for "epistemic mismatch"—where LLMs express unwarranted certainty in their generated outputs despite low internal confidence. Recent empirical studies have shown that models tend to default to confident, assertive language even when their token-level certainty or reasoning reliability is low (Leng et al. (2025); Zhou et al. (2025); Wen et al. (2024)). This growing literature underscores the need for methods that can measure and calibrate linguistic certainty directly, particularly in high-stakes domains.
@@
Studies have attempted to bridge the gap between internal confidence and linguistic assertiveness but still face considerable limitations. Although Mielke et al. (2022) explore epistemic calibration, their study uses a limited scoring scale to measure both assertiveness and confidence, which restricts the continuous assessment of LLM output. Their approach also relies on a narrow range of datasets, thereby limiting its applicability across domains. Zhou et al. (2024a) address miscalibration using epistemic markers, but their method lacks real domain grounding and fails to consider the complexity of natural language.
Framing these developments within the broader context of trust and epistemic communication, concurrent work by Yona et al. (2024) provides strong empirical evidence that large language models often express undue linguistic confidence—answering decisively even when internally uncertain. This misalignment, which they formalize as faithful response uncertainty, reveals that prompting alone is insufficient to ensure models communicate their uncertainty reliably. The resulting “epistemic mismatch” not only hinders trustworthiness but also exacerbates overreliance in knowledge-seeking tasks. Our work contributes to this conversation by offering a model-driven approach to quantifying assertiveness, enabling more precise detection of cases where linguistic style diverges from underlying confidence.
@@
This review of existing work on LLM calibration and confidence reveals several limitations that our research aims to address:
@@
  • Few Integrated Approaches: Previous studies typically address either internal certainty or linguistic assertiveness but rarely both simultaneously (Jiang et al., 2021). There is a need for comprehensive frameworks that integrate these aspects to ensure LLMs communicate accurately and responsibly.
@@
  • Inadequate Assertiveness Measurement: Existing methods for measuring linguistic assertiveness often rely heavily on lexicon-based approaches (Pei & Jurgens, 2021) or subjective perceptions without adequate validation (Steyvers et al., 2024). These methods can lack contextual depth and fail to generalize across diverse domains.
@@
  • Limited High-Stakes Evaluation: Although some studies explore epistemic calibration, they mainly cover a narrow range of topics and employ low-resolution measures of assertiveness, limiting their applicability in critical domains such as misinformation detection (Mielke et al., 2022).
@@
This study also presents a human-centered approach for developing a robust assertiveness scoring method. To ensure reliability, we train and compare several models to identify the best estimator for assertiveness based on accuracy and transferability using a new multi-domain dataset hand-coded to measure certainty. After selecting the top-performing model, we validate the results using a different human-surveyed dataset (Wang, 2017), which was coded independently by different individuals. This comprehensive methodology enables us to thoroughly assess both the objective and subjective aspects of assertiveness in language model explanations.
Our findings reveal that when the model has low internal certainty, it generates explanations that are significantly over-assertive, meaning the language used implies a higher degree of certainty than is warranted by the model’s actual confidence or accuracy. This miscalibration could lead users to misconstrue the model’s judgments as more reliable than they actually are. More precisely, our results confirm a strong correlation between GPT-4o’s assertiveness scores and human perceptions of assertiveness, but a weak correlation between human perceptions and internal certainty, and an even weaker relationship between GPT-4o model assertiveness and internal certainty. Together, these results provide strong evidence of epistemic miscalibration in LLMs. [^1] For the code and datasets used, refer to our GitHub repository at: https://github.com/ComplexData-MILA/epistemic-integrity.
@@ 2.1.2 External Certainty
Mielke et al. (2022) examine how both humans and LLMs interpret and generate expressions of uncertainty. They find that LLMs such as GPT-4 can map uncertainty expressions to numerical equivalents similarly to humans, though they are more prone to biases rooted in their training data. However, most prior research focuses on how LLMs interpret uncertainty expressions rather than how they generate assertive or uncertain language aligned with their internal confidence scores. These developments also align with broader insights from reinforcement learning with human feedback (RLHF), where model behaviors that mirror human-like rhetorical confidence can be implicitly rewarded, regardless of factual accuracy (Leng et al. (2025)). Studies have also found that prompt-level variations can influence perceived certainty (Zhou et al. (2025)), and that task format and cognitive complexity can exacerbate miscalibration (Chhikara (2025)). These findings further validate the need for domain-agnostic tools that assess and align assertiveness with actual model certainty, motivating our focus on assertiveness quantification.
@@ 2.2 Epistemic Calibration
Figure 1 presents two contrasting examples that highlight varying levels of epistemic calibration in LLM outputs. In both cases, the LLM responds to a user query, and we compute two certainty scores: Internal Certainty and External Certainty (through linguistic assertiveness). In the first example, the internal and external certainty scores are closely aligned, demonstrating high epistemic calibration. Here, the model’s linguistic assertiveness accurately reflects its internal confidence, resulting in a trustworthy output. In contrast, the second example reveals low epistemic calibration. Although the internal certainty score is relatively low, the model’s linguistic assertiveness remains high, indicating overconfidence in its response. In the following analysis, we compute internal certainty using the method outlined by Pelrine et al. (2023) and use linguistic assertiveness, derived from our custom model detailed in Section 3—which is validated against human surveyed human ratings—as a proxy for external certainty quantification.
@@ 3.1 Datasets and Models
Followup Findings (March 2025): We test two more models’ performance in measuring assertiveness:
  • Llama-3.2-3B-Instruct fine-tuned with LoRA with rounding (Hu et al., 2021): Fine-tuned using LoRA similar to the Llama-3.2-1B model, but with assertiveness scores rounded to one decimal place in the fine-tuning data as with GPT-4o with rounding.
  • Llama-3.2-11B-Vision-Instruct fine-tuned with LoRA with rounding (Hu et al., 2021): a larger Llama model trained with the same process as the 3B one.
@@ 3.2 Computing internal certainty and linguistic assertiveness
To estimate the internal certainty of the LLM, we use the verbalized confidence method outlined in Pelrine et al. (2023). This method provides a combination of strong performance and simplicity of use compared to other state-of-the-art methods for the misinformation detection domain (Rivera et al., 2024). Specifically, the model is instructed to provide an explanation for each misinformation classification and assign an uncertainty score. The scores are calibrated on a validation set using Platt’s method (Platt et al., 1999). The robustness of this method is further validated in Appendix B.
@@ 4 Results
Figure 2 is obtained from comparing the seven different methods of assertiveness quantification, evaluating on the test set of our composite dataset. We find that GPT-4o fine-tuned with rounding (training on assertiveness scores rounded to one decimal point) achieves the highest accuracy in predicting human-annotated assertiveness scores. The margin of improvement over the approaches from the literature is very large, cutting Mean Squared Error (MSE) by more than half. To validate the transferability of these results across different domains, we conduct an ablation study by training the model on only four of the five datasets, and subsequently testing the model on the excluded one (as opposed to a standard random split in Figure 2). As shown in Table 1, GPT-4o fine-tuned (with training assertiveness scores rounded to one decimal point) originally achieves the lowest average MSE. Thus, GPT-4o appears well-suited to capturing the linguistic nuances that contribute to perceived assertiveness, even in transfer settings. We later found larger Llama models also achieve strong performance, even outperforming GPT-4o. Due to resource constraints, in subsequent sections we present our original results using GPT-4o as our primary tool. However, we note that the Llama models provide promising open-weight alternatives for measuring the assertiveness of generated explanations in the misinformation detection domain.
Figure 2: Comparison of assertiveness evaluation performance across models. The models are evaluated based on their Mean Squared Error (MSE) relative to the test set, which is a subset of our dataset. Among models tested here, the fine-tuned GPT-4o, trained with assertiveness scores rounded to one decimal point, achieved the lowest MSE, which suggests the highest accuracy in predicting assertiveness.
@@
Fine-tuned Llama-3.2-3B-Instruct (Rounded) 1.89 1.53 2.15 1.85 1.33
Fine-tuned Llama-3.2-11B-Vision-Instruct (Rounded) 0.90 0.83 0.92 0.82 0.90
@@
Table 1: A comparison of model performance when training on four datasets and holding out the fifth for the measurements reported here. Numbers report Mean Squared Error (MSE) for each model.
Initial Findings (November 2024): The fine-tuned GPT-4o model, using assertiveness scores rounded to one decimal place, achieves the best overall performance with an average MSE of approximately $1.0781.078$. This result indicates that the model is the most consistent and reliable for assertiveness scoring across various domains.
Followup Findings (March 2025): We add experiments with Fine-tuned Llama-3.2-3B and Llama-3.2-11B, with the same instructions as Fine-tuned GPT- models are added in this analysis with the same instruction as the previous Fine-tuned GPT-4o. Llama-3.2-11B achieves the best overall performance with an average MSE of approximately $0.870.87$. This provides a strong open-weight alternative to GPT-4o.
@@
Figure 3(a) illustrates a comparison between the probability distributions of certainty scores derived from the certainty calibration method proposed by Pelrine et al. (2023) and assertiveness scores from the best-performing model shown in Figure 2 (Fine-tuned GPT, Rounded), applied to the LIAR misinformation dataset. Notably, while the certainty scores exhibit a wide variance, assertiveness scores are more concentrated toward the middle of the distribution. Additionally, Figure 3(b) reveals a low Spearman correlation (0.3) between the two sets of scores, suggesting that there is a significant misalignment between certainty and assertiveness. We provide examples of both epistemically calibrated and uncalibrated explanations with varying levels of assertiveness in Appendix F. Furthermore, in Appendix G, we test whether strong assertions of uncertainty affect calibration, finding that even when controlling for these cases, the model remains heavily skewed towards over-assertiveness.
@@ 7 Discussion and Conclusion
In this work, we discussed the problem of epistemic calibration for LLMs: ensuring that the confidence expressed in a model’s communication aligns with its underlying reliability. We argued that this normative ideal is critical for LLMs to serve as robust and responsible information sources. Through a decomposition of the problem into external and internal certainty, we developed a framework for understanding and evaluating epistemic calibration in LLMs. Using our approach that greatly improves fidelity of assertiveness measurements compared to prior models, our empirical investigation of a state-of-the-art model reveals significant gaps between the model’s internal confidence estimates and the assertiveness of its generated language. This miscalibration poses risks to users, who may be misled by overconfident model outputs.
@@ 10 Author Contributions
Bijean Ghafouri led the human evaluation of assertivity, took the lead writing the second (current) version of the paper, and gave input and feedback throughout the project. Shahrad Mohammadzadeh led the work towards a stronger assertiveness measurement model, and made significant contributions to the direction, writing, and other parts of the project. James Zhou had the idea of assertiveness miscalibration, created the original framing and first version of the paper, and built and ran the initial experiments. Pratheeksha Nair conducted downstream analysis of assertiveness (mis)calibration and contributed to the writing. Jacob-Junqi Tian fine-tuned Llama-3.2-1B and set up the framework for our composite dataset labeling. Hikaru Tsujimura extended Llama fine-tuning experiments to 3B and 11B models. Mayank Goel expanded the assertivity evaluation experiments. Sukanya Krishna contributed to literature review and writing. Reihaneh Rabbany and Jean-François Godbout advised the project, contributing ideas and feedback. Godbout also helped design the human evaluation and contributed substantially to the writing. Kellin Pelrine supervised the project, providing guidance and feedback at all stages.